OpenStudy (anonymous):

The sum of two numbers is 56. One number is 3 times larger than the other. What are the two numbers?

OpenStudy (anonymous):

I don't know the formula you're suppose to use for this.

OpenStudy (saranya):

This is a simultaneous equation problem. x + y = 56 3x = y Rewrite 3x = y as 3x - y = 0 x + y = 56 3x - y = 0 add the equations together 4x = 56 so x = 14 plugging this back into the second equation gives 3(14) = y so y=42 Check: 14 + 42 = 56 3(14) = 42

OpenStudy (anonymous):

Thank you!